ICM 1998, Berlin, Aug. 18–27

Abstracts of Plenary and Invited Lectures

Section:

0. Plenary Lectures

1991 MS Classification: Moser, Jürgen, ETH Zürich, Switzerland Dynamical Systems – Past and Present

This lecture presents, from a personal point of view, several important advances, solved and unsolved problems in dynamical systems, and shows their interconnection with other fields of mathematics.

ICM 1998 Berlin · Plenary and Invited Lectures